# Deep Discriminative to Kernel Density Graph for In- and Out-of-distribution Calibrated Inference

### v7

@@ Abstract
Deep discriminative approaches like random forests and deep neural networks have recently found applications in many important real-world scenarios. However, deploying these learning algorithms in safety-critical applications raises concerns, particularly when it comes to ensuring confidence calibration for both in-distribution and out-of-distribution data points. Many popular methods for in-distribution (ID) calibration, such as isotonic and Platt’s sigmoidal regression, exhibit excellent ID calibration performance. However, these methods are not calibrated for the entire feature space, leading to overconfidence in the case of out-of-distribution (OOD) samples. On the other end of the spectrum, existing out-of-distribution (OOD) calibration methods generally exhibit poor in-distribution (ID) calibration. In this paper, we address ID and OOD calibration problems jointly. We leveraged the fact that deep models, including both random forests and deep-nets, learn internal representations which are unions of polytopes with affine activation functions to conceptualize them both as partitioning rules of the feature space. We replace the affine function in each polytope populated by the training data with a Gaussian kernel. Our experiments on both tabular and vision benchmarks show that the proposed approaches obtain well-calibrated posteriors while mostly preserving or improving the classification accuracy of the original algorithm for in-distribution region, and extrapolate beyond the training data to handle out-of-distribution inputs appropriately.
${}^{1}$ohns Hopkins University (JHU), ${}^{2}$cale AI $*{}^{*}$ corresponding author: jdey4@jhu.edu, $\dagger$ denotes equal contribution
@@ 1 Introduction
Machine learning methods, specially deep neural networks and random forests have shown excellent performance in many real-world tasks, including drug discovery, autonomous driving and clinical surgery. However, calibrating confidence over the whole feature space for these approaches remains a key challenge in the field. Although these learning algorithms can achieve near optimal performance at inferring on the samples lying in the high density regions of the training data [1, 2, 3], they yield highly confident predictions for the samples lying far away from the training data [4]. Calibrated confidence within the training or in-distribution (ID) region as well as in the out-of-distribution (OOD) region is crucial for safety critical applications like autonomous driving and computer-assisted surgery, where any aberrant reading should be detected and taken care of immediately [4, 5]. A well-calibrated model capable of quantifying the uncertainty associated with inference for any points from the training distribution as well as detecting OOD data can be a life-saver in these cases.
@@
The approaches to calibrate OOD confidence for learning algorithms described in the literature can be roughly divided into two groups: discriminative and generative. Intuitively, the easiest solution for OOD confidence calibration is to learn a function that gives higher scores for in-distribution samples and lower scores for OOD samples, and thereby re-scale the posterior or confidence score from the original model accordingly [6]. The discriminative approaches try to either modify the loss function [7, 8, 9] or train the network exhaustively on OOD datasets to calibrate on OOD samples [10, 4]. Recently, as shown by Hein et al. [4], the ReLU networks produce arbitrarily high confidence as the inference point moves far away from the training data. Therefore, calibrating ReLU networks for the whole OOD region is not possible without fundamentally changing the network architecture. As a result, all of the aforementioned algorithms are unable to provide any guarantee about the performance of the network throughout the whole feature space. The other group tries to learn generative models for the in-distribution as well as the out-of-distribution samples. The general idea is to conduct likelihood ratio tests for a particular sample using the generative models [11], or threshold the ID likelihoods to detect OOD samples. However, it is not obvious how to control likelihoods far away from the training data for powerful generative models like variational autoencoders (VAEs) [12] and generative adversarial networks (GAN) [13]. Moreover, Nalisnick et al. [14] and Hendrycks et al. [10] showed VAEs and GANs can also yield overconfident likelihoods far away from the training data.
The algorithms described so far are concerned with OOD confidence calibration for deep-nets only. However, in this paper, we show that other approaches which partition the feature space, for example random forest, can also suffer from poor confidence calibration both in the ID and the OOD regions. Moreover, the algorithms described above are concerned about the confidence in the OOD region only and do not address the confidence calibration within the in-distribution at all. This issue is addressed separately in a different group of literature [15, 16, 17, 18, 19, 20]. Instead, we consider both calibration problems jointly and propose an approach that achieves good calibration throughout the whole feature space.
In this paper, we conceptualize both random forest and ReLU networks as partitioning rules with an affine activation over each polytope. We consider replacing the affine functions learned over the polytopes with Gaussian kernels. We propose two novel kernel density estimation techniques named Kernel Density Forest (KDF) and Kernel Density Network (KDN). Our proposed approach completely excludes the need for training on OOD examples for the model (unsupervised OOD calibration). We conduct several simulation and real data studies that show both KDF  and KDN  are well-calibrated for OOD samples while they maintain good performance in the ID region.
2 Related Works and Our Contributions
There are a number of approaches in the literature which attempt to learn a generative model and control the likelihoods far away from the training data. For example, Ren et al. [11] employed likelihood ratio test for detecting OOD samples. Wan et al. [8] modified the training loss so that the downstream projected features follow a Gaussian distribution. However, there is no guarantee of performance for OOD detection for the above methods. To the best of our knowledge, apart from us, only Meinke et al. [5] has proposed an approach to guarantee asymptotic performance for OOD detection. They model the training and the OOD distribution using Gaussian mixture models which enable them to control the class conditional posteriors far away from the training points. Compared to the aforementioned methods, our approach differs in several ways:
  • We address the confidence calibration problems for both ReLU-nets and random forests from a common ground.
  • We address in-distribution (ID) and out-of-distribution (OOD) calibration problem as a continuum rather than two separate problems.
  • We provide an algorithm for OOD confidence calibration for both tabular and vision datatsets whereas most of the OOD calibration methods in the literature are tailor-made for vision problems.
  • We propose an unsupervised OOD calibration approach, i.e., we do not need to train exhaustively on different OOD samples.
3 Technical Background
@@ 3.1 Setting
Consider a supervised learning problem with independent and identically distributed training samples $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ such that $(\mathbf{X},Y)\sim P_{X,Y}$, where $\mathbf{X}\sim P_{X}$ is a $\mathcal{X}\subseteq\mathbb{R}^{D}$ valued input and $Y\sim P_{Y}$ is a $\mathcal{Y}=\{1,\cdots,K\}$ valued class label. Let $\mathcal{S}$ be a high density region of the marginal, $P_{X}$, thus $\mathcal{S}\subsetneq\mathcal{X}$
Definition 3.1 (In- and Out-of-distribution Samples).
If $\mathbf{x}\in\mathcal{S}$, it is an in-distribution (ID) sample and if $\mathbf{x}\notin\mathcal{S}$, it is an out-of-distribution (OOD) sample.
Here the goal is to learn a confidence score, $\mathbf{g}:\mathbb{R}^{D}\rightarrow[0,1]^{K}$, $\mathbf{g}(\mathbf{x})=[g_{1}(\mathbf{x}),g_{2}(\mathbf{x}),\dots,g_{K}(\mathbf{x})]$ such that,
$g_{y}(\mathbf{x})=\begin{cases}P_{Y|X}(y|\mathbf{x}),&\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \mathbf{x}\in\mathcal{S}\\ P_{Y}(y),&\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \mathbf{x}\notin\mathcal{S}\end{cases},\quad\forall y\in\mathcal{Y}$ (1)
where $P_{Y|X}(y|\mathbf{x})$ is the posterior probability for class $y$ given by the Bayes formula:
@@
Here $P_{X|Y}(\mathbf{x}|y)$ is the class conditional density for the training data which we will refer as $f_{y}(\mathbf{x})$ hereafter for brevity.
@@ 3.2 Main Idea
Deep discriminative networks partition the feature space $\mathbb{R}^{d}$ into a union of $p$ affine polytopes $Q_{r}$ such that $\bigcup_{r=1}^{p}Q_{r}=\mathbb{R}^{d}$, and learn an affine function over each polytope [4, 21]. Mathematically, the class-conditional density for the label $y$ estimated by these deep discriminative models at a particular point $\mathbf{x}$ can be expressed as:
@@
For example, in the case of a decision tree, $\mathbf{a}_{r}=\mathbf{0}$, i.e., decision tree assumes uniform distribution for the class-conditional densities over the leaf nodes. Among these polytopes, the ones that lie on the boundary of the training data extend to the whole feature space and hence encompass all the OOD samples. Since the posterior probability for a class is determined by the affine activation over each of these polytopes, the algorithms tend to be overconfident when making predictions on the OOD inputs. Moreover, there exist some polytopes that are not populated with training data. These unpopulated polytopes serve to interpolate between the training sample points. If we replace the affine activation function of the populated polytopes with Gaussian kernels and prune the unpopulated ones, the tail of the kernel will help interpolate between the training sample points while assigning lower likelihood to the low density or unpopulated polytope regions of the feature space. This results in better confidence calibration for the proposed modified approach.
@@ 3.3 Proposed Approach
We will call the above discriminative approaches as the ‘parent approach’ hereafter. Consider the collection of polytope indices $\mathcal{P}$ from the parent approach which contains the indices of total $\tilde{p}$ polytopes populated by the training data. We replace the affine functions over the populated polytopes with Gaussian kernels $\mathcal{G}(\cdot;\hat{\mu}_{r},\hat{\Sigma}_{r})$. For a particular inference point $\mathbf{x}$, we consider the Gaussian kernel with the minimum distance from the center of the kernel to the corresponding point:
$r^{*}_{\mathbf{x}}=\operatornamewithlimits{argmin}_{r}\|\mu_{r}-\mathbf{x}\|,$ (4)
[…]
$\tilde{f}_{y}(\mathbf{x})=\frac{1}{n_{y}}\sum_{r\in\mathcal{P}}n_{ry}\mathcal{G}(\mathbf{x};\mu_{r},\Sigma_{r})\mathbbm{1}(r=r^{*}_{\mathbf{x}}),$ (5)
where $n_{y}$ is the total number of samples with label $y$ and $n_{ry}$ is the number of samples from class $y$ that end up in polytope $Q_{r}$. We add a small constant to the class conditional density $\tilde{f}_{y}$:
@@
Note that in Equation 6, $\frac{b}{\log(n)}\rightarrow 0$ as the total training points, $n\rightarrow\infty$. The intuition behind the added constant will be clarified further later in Proposition 4.4. The confidence score $\hat{g}_{y}(\mathbf{x})$ for class $y$ given a test point $\mathbf{x}$ is estimated using the Bayes rule as:
@@
where $\hat{P}_{Y}(y)$ is the empirical prior probability of class $y$ estimated from the training data. We estimate the class for a particular inference point $\mathbf{x}$ as:
@@
3.4 Desiderata
We desire our proposed model to estimate confidence score $\hat{g}_{y}$ to satisfy the following three desiderata:
  1. 1.
    Empirical Classification Accuracy : For finite sample cases, our goal is to maintain or even enhance the parent approach’s ID accuracy on standard benchmark datasets.
  2. 2.
    Empirical ID Calibration On standard benchmark datasets, we aim to enhance the ID calibration of the parent approach.
  3. 3.
    Empirical OOD Calibration We aim to enhance the OOD calibration of the parent appraoch against any OOD datasets.
We run extensive experiments on various simulation and benchmark datasets in Section 5 to empirically verify that our proposed approach achieves Desiderata 1, 2 and 3.
4 Kernel Density Graph
In this section we describe how we estimate the parameters shown in Equation 5.

### v9

# Simple Calibration via Geodesic Kernels

@@ Abstract
Deep discriminative approaches, such as decision forests and deep neural networks, have recently found applications in many important real-world scenarios. However, deploying these learning algorithms in safety-critical applications raises concerns, particularly when it comes to ensuring calibration for both in-distribution and out-of-distribution regions. Many popular methods for in-distribution (ID) calibration, such as isotonic and Platt’s sigmoidal regression, exhibit adequate ID calibration performance. However, these methods are not calibrated for the entire feature space, leading to overconfidence in the out-of-distribution (OOD) region. Existing OOD calibration methods generally exhibit poor ID calibration. In this paper, we jointly address the ID and OOD problems. We leveraged the fact that deep models learn to partition feature space into a union of polytopes, that is, flat-sided geometric objects. We introduce a geodesic distance to measure the distance between these polytopes and further distinguish samples within the same polytope using a Gaussian kernel. Our experiments on both tabular and vision benchmarks show that the proposed approaches, namely Kernel Density Forest (KDF) and Kernel Density Network (KDN), obtain well-calibrated posteriors for both ID and OOD samples, while mostly preserving the classification accuracy and extrapolating beyond the training data to handle OOD inputs appropriately.
@@ 1 Introduction
Machine learning methods, specifically deep neural networks and decision forests (such as random forests and gradient boosting trees), show excellent performance in many real-world tasks (Xu et al., 2021a), including drug discovery (Cano et al., 2017; Shi et al., 2019), autonomous driving (Yurtsever et al., 2020), and clinical surgery (Bihorac et al., 2019). However, Calibrating confidence across the entire feature space for these approaches continues to be a significant challenge (Hein et al., 2019), as most existing algorithms focus either on in-distribution (ID) calibration or out-of-distribution (OOD) detection, rather than treating them as a unified problem. Calibrated confidence within the training or in-distribution (ID) region as well as in the out-of-distribution (OOD) region is crucial for safety critical applications like autonomous driving and computer-assisted surgery, where any aberrant reading should be detected and taken care of immediately (Hein et al., 2019; Meinke et al., 2021).
@@
The methods for detecting out-of-distribution (OOD) samples in the literature can be broadly categorized into two types: discriminative and generative approaches. A straightforward strategy for OOD detection is to design a function that assigns higher scores to in-distribution (ID) samples and lower scores to OOD samples (Liang et al., 2017). Discriminative approaches typically modify the loss function (Nandy et al., 2020; Wan et al., 2018; DeVries & Taylor, 2018) or extensively train the model on OOD datasets to produce these distinct scores (Hendrycks et al., 2018; Hein et al., 2019). Recently, Hein et al. (2019) showed ReLU networks produce arbitrarily high confidence as the inference point moves far away from the training data and it prohibits most algorithms from having an asymptotic performance guarantee. Therefore, calibrating ReLU networks for the whole OOD region is not possible without fundamentally changing the network architecture. Others learn generative models for the ID and the OOD samples. The general idea is to perform the likelihood ratio test for a particular sample using generative models Ren et al. (2019), or to threshold the ID likelihoods to detect OOD samples. However, it is not obvious how to control the likelihoods far away from the training data for powerful generative models such as variational autoencoders (VAE) (Kingma et al., 2019) and adversarial generative networks (GAN) (Goodfellow et al., 2020). Moreover, Nalisnick et al. (2018) and Hendrycks et al. (2018) showed that VAEs and GANs can yield overconfident likelihoods far away from the training data. Although these algorithms focus solely on OOD detection for deep networks, other methods, such as those that partition the feature space (e.g. decision forests), can also exhibit overconfidence in OOD regions.
The algorithms described above are concerned only with OOD detection and do not address confidence calibration within the ID region at all. This is addressed by a separate group of algorithms (Zadrozny & Elkan, 2001; Caruana, 2004; Platt et al., 1999; Guo et al., 2017; 2019; Kull et al., 2019) that deals with ID confidence calibration.
We estimate posteriors that are calibrated in both ID and OOD regions. Our approach subsumes OOD detection and ID calibration by addressing the calibration problem as a continuum from ID to OOD regions (see Appendix Section A for further clarification). We conceptualize decision forests and ReLU networks as partitioning the input space into unions of polytopes, where each polytope has a distinct affine transformation. Nearby polytopes have similar affine functions. We used a “Geodesic” kernel to measure distances between these polytopes (Madhyastha et al., 2020), allowing us to better capture their relationships. We find the nearest polytope to an inference point and replace the affine function learned over that polytope with a Gaussian kernel, leading to two novel calibration methods: Kernel Density Forest (KDF) and Kernel Density Network (KDN). Our approach eliminates the need for training on OOD examples, enabling fully unsupervised OOD calibration. Through extensive simulations and benchmark experiments, we show that KDF and KDN achieve well-calibrated predictions for OOD samples while maintaining good accuracy and calibration in the ID region, on both tabular and vision benchmarks.
2 Methods
@@ 2.1 Setting
Consider a supervised learning problem with independent and identically distributed training samples $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ such that $(\mathbf{X},Y)\sim P_{X,Y}$, where $\mathbf{X}\sim P_{X}$ is a $\mathcal{X}\subseteq\mathbb{R}^{D}$ valued input and $Y\sim P_{Y}$ is a $\mathcal{Y}=\{1,\cdots,K\}$ valued class label. Let $\mathcal{S}$ be the high density region of the marginal, $P_{X}$, thus $\mathcal{S}\subsetneq\mathcal{X}$. The posterior probability for class $y$ is given by the Bayes formula:
@@
Here $P_{X|Y}(\mathbf{x}|y)$ is the class conditional density. In the OOD region where $\mathbf{x}\notin\mathcal{S}$, we do not have any evidence to favor a particular class $y$ for an inference point $\mathbf{x}$. Hence we can write:
$P_{X|Y}(\mathbf{x}|y_{i})=P_{X|Y}(\mathbf{x}|y_{j}),\forall y_{i},y_{j}\in\mathcal{Y}.$ (2)
We will refer to $P_{X|Y}(\mathbf{x}|y)$ as $f_{y}(\mathbf{x})$ hereafter for brevity. Substituting Equation 2 into Equation 1 for a data point in the out-of-distribution region yields:
$P_{Y|X}(y|\mathbf{x})=\frac{P_{Y}(y)}{\sum_{k=1}^{K}P_{Y}(k)}=P_{Y}(y).$ (3)
Our goal is to learn a confidence score, $\mathbf{g}:\mathbb{R}^{D}\rightarrow[0,1]^{K}$, $\mathbf{g}(\mathbf{x})=[g_{1}(\mathbf{x}),g_{2}(\mathbf{x}),\dots,g_{K}(\mathbf{x})]$ such that,
$g_{y}(\mathbf{x})=\begin{cases}P_{Y|X}(y|\mathbf{x}),&~{}\text{if}~{}\mathbf{x}\in\mathcal{S}\\ P_{Y}(y),&~{}\text{if}~{}\mathbf{x}\notin\mathcal{S}\end{cases},\quad\forall y\in\mathcal{Y}$ (4)
@@ 2.2 Main Idea
Deep discriminative networks partition the feature space $\mathbb{R}^{d}$ into a union of $p$ affine polytopes $Q_{r}$ such that $\bigcup_{r=1}^{p}Q_{r}=\mathbb{R}^{d}$, and learn an affine function over each polytope (Hein et al., 2019; Xu et al., 2021b; Black et al., 2022; Priebe et al., 2020). Mathematically, the unnormalized class-conditional density for the label $y$ estimated by these deep discriminative models at a particular point $\mathbf{x}$ can be expressed as:
@@
For example, in the case of a decision tree, $\mathbf{a}_{r}=\mathbf{0}$, i.e., decision trees assume a uniform distribution for the class-conditional densities within a leaf node. Among these polytopes, the ones that lie on the boundary of the training data extend to the whole feature space and hence encompass all over the OOD region. Since the posterior probability for a class is determined by the affine activation over each of these polytopes, the algorithms tend to be overconfident when making predictions on the OOD inputs. Moreover, there exist some polytopes that are not populated with training data. These unpopulated polytopes serve to interpolate between the training sample points. If we replace the affine activation function of the polytope with a Gaussian kernel and prune the polytopes unpopulated by the training data, the tail of the kernel will help interpolate between the training sample points while assigning lower likelihood to the low density or unpopulated polytope regions of the feature space.
@@ 2.3 Proposed Approach
We will call the above discriminative approaches as the ‘parent approach’ hereafter. Consider the collection of polytope indices $\mathcal{P}$ from the parent approach which are populated by the training data. We replace the affine functions over the populated polytopes with Gaussian kernels $\mathcal{G}(\cdot;\hat{\mu}_{r},\hat{\Sigma}_{r})$. For a particular inference point $\mathbf{x}$, we consider the Gaussian kernel with the minimum distance from the center of the kernel to the corresponding point:
$\hat{r}^{*}_{\mathbf{x}}=\operatornamewithlimits{argmin}_{r}\|\mu_{r}-\mathbf{x}\|,$ (6)
where $\|x_{1}-x_{2}\|$ denotes a distance between $x_{1}$ and $x_{2}$. As we will show later, the type of distance considered in Equation 6 highly impacts the performance of the proposed model. In short, we modify Equation 5 from the parent ReLU-net or decision forest to estimate the class-conditional density:
$\tilde{f}_{y}(\mathbf{x})=\frac{1}{n_{y}}\sum_{r\in\mathcal{P}}n_{ry}\mathcal{G}(\mathbf{x};\hat{\mu}_{r},\hat{\Sigma}_{r})\mathbbm{1}(r=\hat{r}^{*}_{\mathbf{x}}),$ (7)
where $n_{y}$ is the total number of training samples with label $y$ and $n_{ry}$ is the number of training samples from class $y$ that end up in polytope $Q_{r}$. Note that our proposed model is not a Gaussian mixture like model; rather, for a given inference point, it considers only the single nearest Gaussian. We add a small constant to the class conditional density so that the class conditional likelihoods become constant far away from the Gaussian center:
@@
Note that in Equation 8, $\frac{b}{\log(n)}\rightarrow 0$ as the total training points, $n\rightarrow\infty$. See the derivation of Proposition 2 in Appendix B for further clarification on how the above constant helps in achieving our goal in Equation 4. The confidence score $\hat{g}_{y}(\mathbf{x})$ for class $y$ given a test point $\mathbf{x}$ is estimated using the Bayes rule as:
@@
where $\hat{P}_{Y}(y)$ is the empirical prior probability of class $y$ estimated from the training data. We obtain the class for a particular inference point $\mathbf{x}$ by:
@@
2.4 Geodesic Kernel
Figure 1: Geodesic kernel calculation for two samples in decision forest and ReLU-net. Two samples $\mathbf{x}_{1}$ (marked blue) and $\mathbf{x}_{2}$ (marked red) are pushed down A) a decision forest with two trees. B) a ReLU-net with three layers. The ReLU activation for the two samples at each node is shown with colors blue and red. The activation paths that are identically activated for the two samples above are bolded, along with the common elements in their activation modes.
As evident in Equation 5, samples that fall within the same polytope undergo the same affine transformation by the deep network. Black et al. (2022) identified polytopes as the fundamental building blocks of deep networks, introducing the concept of a "spline code," which represents the binarized and flattened activation pattern of the network for an input sample. These spline codes uniquely identify the polytope in which a sample falls and, as demonstrated by Black et al. (2022), can effectively cluster identical samples without even considering the absolute magnitude of activations. They further proposed using the Hamming distance between spline codes to identify nearby polytopes undergoing nearly similar transformations to a certain polytope in the network. Moreover, various other kernels based on activation patterns in deep networks have been explored in the literature (Balestriero et al., 2018; Wilson et al., 2016). However, to our knowledge, most of these kernels either do not construct a metric space or at most construct a semi-metric space and hence, they cannot be used as valid metrics (Burago, 2001). In the following, we propose a generalized kernel for both the decision forest and ReLU-net, and, as we will show in Proposition 1, it constructs a metric space. This enables us to use our kernel to measure the distance between samples in the representation space learned by the network.
Our proposed kernels are based on existing random partition kernels (Davies & Ghahramani, 2014). Hence, all the properties of such partition kernels hold for our proposed kernel. The only difference is that—- we do not use random partitions of the feature space, instead we use the partitions learned by decision forests or ReLU-net after being trained on the training data. We identify these partitions using the leaf identities in decision forest or the activation patterns in a ReLU-net. Note that the geometry of these partitions or polytopes encodes the low-dimensional structure learned by the model (Balestriero et al., 2018).
2.4.1 Forest Kernel
A decision forest with $B$ decision trees partitions the feature space into a set of disjoint polytopes, $\{Q_{r}\}_{r=1}^{p}$. Decision forests map any sample $\boldsymbol{x}_{i}\in Q_{r}$ to a $B$ dimensional vector $\boldsymbol{a}_{i}$ where each element $a_{i}(t)$ of the vector represents the leaf identities the sample falls in each tree. We call the above vector $\boldsymbol{a}_{i}$ the “activation mode”.
If the two activation modes for two different training samples are identical, they belong to the same polytope. In other words, if $\boldsymbol{a}_{i}=\boldsymbol{a}_{j}$, then $Q_{r}=Q_{s}$. This statement holds because the above samples will end up in the same partition in all decision trees.
For any two points $\mathbf{x}_{i}\in Q_{r}$ and $\mathbf{x}_{j}\in Q_{s}$, we define the kernel $\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ as:
$\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})=\frac{1}{B}\sum_{t=1}^{B}\mathbbm{1}(a_{i}(t)=a_{j}(t)).$ (11)
Note that $0\leq\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})\leq 1$. In short, $\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ is the fraction of total trees where the two samples follow the same path from the root to a leaf node.
2.4.2 Network Kernel
A fully connected $L$ layer ReLU-net partitions the feature space into a set of disjoint polytopes, $\{Q_{r}\}_{r=1}^{p}$. We have the set of all nodes in a particular layer $l$ denoted by $\mathcal{N}_{l}=\{n_{l}^{i}\}_{i=1}^{|\mathcal{N}_{l}|}$. We can randomly pick a node $n_{l}^{i_{l}}\in\mathcal{N}_{l}$ at each layer $l$, and construct a sequence of nodes starting at the input layer and ending at the output layer, which we call an activation path: $m=\{n_{l}^{i_{l}}\in\mathcal{N}_{l}\}_{l=1}^{L}$. Note that there are $B=\Pi_{l=1}^{L}{|\mathcal{N}_{l}|}$ possible activation paths for a sample in the ReLU-net. We index each path by a unique identifier number $t\in\{1,\cdots,B\}$.
While pushing a training sample $\mathbf{x}_{i}\in Q_{r}$ through the network, we define the activation from a ReLU unit at any node as ‘$1$’ when it has positive output and ‘00’ otherwise. Therefore, the activation indicates on which side of the affine function at each node the sample falls. The activation of all nodes on an activation path $m_{t}$ for a particular sample creates a string of binary numbers $a_{i}(t)\in\{0,1\}^{L}$. Therefore, a ReLU-net maps any sample $\boldsymbol{x}_{i}\in Q_{r}$ to a $B$ dimensional vector $\boldsymbol{a}_{i}$ where each element $a_{i}(t)$ of the vector represents a string of binary numbers of length $L$. We call the above vector $\boldsymbol{a}_{i}$ the “activation mode”. Note that similar to the leaf identities in decision forests, each element $a_{i}(t)$ of the above vector represents a partition of the feature space, that is, only the samples belonging to a particular region of the feature space can have $a_{i}(t)$. The intersection of all such overlapping partitions in an activation mode gives the polytope $Q_{r}$ in which the sample falls.
If the two activation modes for two different training samples are identical, they belong to the same polytope. In other words, if $\boldsymbol{a}_{i}=\boldsymbol{a}_{j}$, then $Q_{r}=Q_{s}$. This statement holds because the above samples will lie on the same side of the affine function at each node in different layers of the network. Now, we define the kernel $\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ as:
$\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})=\frac{1}{B}\sum_{t=1}^{B}\mathbbm{1}(a_{i}(t)=a_{j}(t)).$ (12)
Note that $0\leq\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})\leq 1$. In short, $\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ is the fraction of total activation paths that are activated identically for two samples in two different polytopes $r$ and $s$. A generalized pseudocode outlining the above two kernel calculation is provided in Algorithm 1. See Figure 1 for a visual interpretation of the kernel calculation.
Algorithm 1 Computing Geodesic Kernel
1:
2:(1) $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbb{R}^{1\times d}$triangleright two input samples
3:(2) $\theta$triangleright parameters for parent decision forest or ReLU-net
4:$\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})\in[0,1]$triangleright compute similarity between $i$ and $j$-th samples.
5:function computeKernel($\mathbf{x}_{i},\mathbf{x}_{j},\theta$)
6:     $\boldsymbol{a}_{i}\leftarrow$ pushDownNetwork($\mathbf{x}_{i},\theta$) $\triangleright$ push samples down the deep network to get the activation modes.
7:     $\boldsymbol{a}_{j}\leftarrow$ pushDownNetwork($\mathbf{x}_{j},\theta$)
8:     $M\leftarrow$ countMatches($\boldsymbol{a}_{i},\boldsymbol{a}_{j}$) $\triangleright$ count the number of similar activation modes or leaf identities
9:     $\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})\leftarrow\frac{M}{B}$triangleright $B$ is the dimension of activation modes.
10:     return $\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$
11:end function
2.5 Network Kernel with Deep Network
Figure 2: Average kernel values are higher for similar-class pairs than for dissimilar ones. Values decrease with depth in networks with random partition (left) but remain stable when trained on data (right).
In this section, we empirically assess the validity of the kernel defined in Equation 12 as network depth increases. We conduct experiments on a two-dimensional Gaussian XOR simulation (described in Appendix D) and the 784-dimensional Fashion-MNIST dataset (from OpenML-CC18 (Bischl et al., 2017)) using fully-connected networks of varying depth and width. Specifically, we measure the average kernel similarity between sample pairs from the same and different class labels, both before training (with random initialization) and after training. As shown in Figure 2, kernel similarity decreases monotonically with depth under random initialization, indicating increasingly disjoint activation paths. In contrast, after training, similarity between same-class samples remains relatively high, suggesting that learning encourages alignment in activation paths. These results demonstrate that, while the kernel becomes sparse in untrained deep networks with random weights, it retains its discriminative power once the network is trained.
2.6 Geodesic Distance
Consider $\mathcal{P}_{n}=\{Q_{1},Q_{2},\cdots,Q_{p}\}$ as a partition of $\mathbb{R}^{d}$ given by a decision forest or a ReLU-net after being trained on $n$ training samples. We measure distance between two points $\mathbf{x}_{i}\in Q_{r},\mathbf{x}_{j}\in Q_{s}$ using the kernel introduced in Equation 11 and Equation 12, and call it ‘Geodesic’ distance (Schölkopf, 2000):
$d(\boldsymbol{x}_{i},\boldsymbol{x}_{j})=-\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})+\frac{1}{2}(\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{i})+\mathcal{K}(\boldsymbol{x}_{j},\boldsymbol{x}_{j}))=1-\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j}).$ (13)
Proposition 1.
$(\mathcal{P}_{n},d)$ is a metric space.
Proof.
See Appendix B.1 for the proof. ∎
The above proposition enables us to use Geodesic distance as a metric in our algorithm. We use Geodesic distance in Equation 6 to find the nearest polytope to the inference point. As Geodesic distance cannot distinguish between points within the same polytope, it has a resolution similar to the size of the polytope. We assume the polytope is small enough so that the manifold within a polytope can be considered Euclidean and hence for discriminating between two points within the same polytope, we fit a Gaussian kernel within the polytope (described below). Moreover, Gaussian kernels decay exponentially from their center which improves our OOD calibration. In Section 3, we will empirically show that using Geodesic distance scales better with higher dimensions compared to Euclidean distance.